# In Search of a Hidden Curve

### v1

@@ Abstract.
Three enumeration problems are discussed in these lectures. One is a detailed account on simple Hurwitz numbers, explaining how the problem was solved by discovering a key curve. This key ignites the generating functions of Hurwitz numbers and drives them into polynomials. The unexpected polynomiality then brings us simple proofs of the Witten-Kontsevich theorem and the $\lambda_{g}$-theorem of Faber-Pandharipande. An analogous problem associated with Catalan numbers is also presented, which has a simpler feature in terms of analysis. The third enumeration problem is a quest of identifying a hidden curve behind the Apéry sequence, which remains to be discovered.
It is examined how a counting problem leads to a spectral curve through a differential equation, of Picard-Fuchs type but can be nonlinear. The counting problem is a geometric question associated with the genus 00, one marked point case. Going from the $(0,1)$-case to arbitrary $(g,n)$-case is explained as a process of quantization of the spectral curve.
This perspective of quantization is discussed in a geometric setting, when the differential equations have holomorphic coefficients, in terms of Higgs bundles, opers, and Gaiotto’s conformal limit construction. In this context, however, there are no counting problems behind the scene.
@@ 0.1. The story begins with Catalan numbers and their Laplace transform
as a basis for all solutions. It then reminds us of the quadratic formula for $z^{2}-xz+1=0$, and a possible relation to the second equation through the Weyl quantization $\begin{cases}z\longmapsto-\hbar\frac{d}{dx},\\ x\longmapsto x.\end{cases}$ Equivalently, we can find the same polynomial from the semi-classical limit of (0.2):
$\lim_{\hbar\rightarrow 0}e^{-\frac{S_{0}(x)}{\hbar}}\left(\hbar^{2}\frac{d^{2}}{dx^{2}}+\hbar x\frac{d}{dx}+1\right)e^{\frac{S_{0}(x)}{\hbar}}=\big{(}S_{0}(x)^{\prime}\big{)}^{2}+xS_{0}(x)^{\prime}+1,\quad z=-S_{0}(x)^{\prime}.$
@@
The curve in common appears to be simply $x=z+\frac{1}{z}.$ Is there anything significant here? Indeed, this curve will tell us a rich story of geometry that is not obvious at all from the shape of these differential equations. As we see in the main text (1.2), $z(x)$ is an unconventional generating function of Catalan numbers $C_{m}:1,2,5,14,42,132,429,1430,4862,16796\dots$ One interpretation of these numbers is the count of cell-decompositions of a two-dimensional sphere $S^{2}$ with one 00-cell and $m$ $1$-cells, while rotation symmetry is not allowed to avoid complication coming from automorphism count (see Remark 2.29).
Figure 0.2. A cell-decomposition of a curve of genus 00 with one 00-cell. The arrow on a half-edge indicates no rotation symmetry is allowed.
@@
What do these geometric quantities of moduli spaces have anything to do with the differential equation (0.2) above? We will see that its solution $\psi(x,\hbar)$ is the generating function of these generating functions $F_{g,n}$ for all $g\geq 0$ and $n>0$. In this context, since the curve $x=z+1/z$ is the semi-classical limit of (0.2), the differential equation is called the quantum curve [1, 22, 50] associated with the classical curve $x=z+1/z$. The Catalan numbers are the count of cell-decompositions of the $(g,n)=(0,1)$ geometry. Hence going from $z(x)$ to $F_{g,n}(t_{1},\dots,t_{n})$ for arbitrary $(g,n)$ is the process of quantization.
At the same time, the fact that generating functions $F_{g,n}(t_{1},\dots,t_{n})$ know the $\psi$-class intersection numbers of ¯${\overline{\mathcal{M}}}_{g,n}$ indicates that they give the Gromov-Witten invariants $GW_{g,n}(\bullet)$ of a point. Therefore, from the point of view of Mariño [65] and Bouchard-Klemm-Mariño-Pasquetti [13], we understand that the spectral curve $x=z+1/z$ is the mirror $B$-model, corresponding to the Gromov-Witten invariants on the $A$-model. From the Catalan numbers to this point of view of mirror symmetry is the theme of Section 1.1.
The polynomiality of some quantities appearing in enumerative problems associated with the moduli spaces ¯${\overline{\mathcal{M}}}_{g,n}$ was discovered in the author’s earlier paper [76], which was the key to the solution [36] of the Hurwitz number conjecture of Bouchard and Mariño [14]. The polynomiality of the generating functions of simple Hurwitz numbers established in [76] presented another surprise: it gives simple few-line proofs of Witten’s conjecture [95] and the $\lambda_{g}$-conjecture of Faber-Pandharipande [38, 39].
The stories coming out of Hurwitz numbers are weaved in Section 2. The hidden curve in this context is the Lambert curve (2.15), whose role in discovering the polynomiality is explained in detail. Bouchard and Mariño [14] identified the Lambert curve through the limiting process of the mirror curves of toric Calabi-Yau $3$-folds [13], and using Lagrange’s Inversion Formula [94]. Appendix gives an elementary analysis that leads to this identification of the Lambert curve. In Section 2.6, we will show that the Lambert curve is actually a simple consequence of tree counting, and the formula for the curve, i.e., the Lambert function, directly follows from a combinatorial identity of trees.
@@
    $z(x)=\sum_{m=0}^{\infty}\frac{C_{m}}{x^{2m+1}},\qquad x=z+\frac{1}{z}$
@@
  • This effect is explained in detail in Section 2 using Hurwitz numbers. The starting sequence is the number of rooted trees on $k$ nodes and their generating function,
@@
  • In Section 2.3 it is explained that the Laplace transform of the $(g,n)$-Hurwitz number in the stable region $02g-2+n>$ 0 is a polynomial in the $t$-variables. We will prove that the Laplace transform of the combinatorial formula, the cut-and-join equation of [46, 91], is a differential equation, and that it automatically proves the Witten-Kontsevich theorem on the $\psi$-class intersection numbers and the Faber-Pandharipande theorem on the $\lambda_{g}$-conjecture [76].
@@ 0.2. A miraculous integer sequence
$\widetilde{GW}_{0,1}(n):=(-1)^{n}(n!)^{2}\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}\bigg{(}1+(m-\ell)\big{(}H_{2\ell+m}+2H_{m+2\ell}-5H_{m}\big{)}\bigg{)}$
@@
Then we have (see [45, 96]):
(0.3) $A_{n}=\widetilde{GW}_{0,1}(n)\qquad\text{for all }\;n.$
The first few terms are: $1,5,73,1445,33001,819005,21460825,584307365,16367912425,\dots$. What is unbelievable about this equality is that it is not obvious at all that $\widetilde{GW}_{0,1}(n)$ as defined by the formula gives a positive integer sequence, let alone it is equal to the second line. What we know about (0.3) so far includes the following:
  • The formula comes from the mirror symmetry, a geometric relation between the Gromov-Witten invariants of a $3$-dimensional variety and the Gauss-Manin connection associated with its mirror partner. In this particular context, it is proved in Golyshev-Zagier [45]. We refer to Zagier [96] for an inspiring account of this interplay.
  • The quantity $\widetilde{GW}_{0,1}(n)$, the quantum period, was obtained by Coates, Corti, Galkin, and Kasprzyk [18], which is essentially the same as the genus 00, $1$-marked point case of the degree $n$ Gromov-Witten invariant of a Fano $3$-fold known as $V_{12}$. The generating function of these numbers, after appropriate modifications including the Borel-Laplace transform, satisfies a Quantum Differential Equation. A brief background will be explained in Section 4.
@@ 0.3. The unique quantization for the holomorphic cases
It is therefore rather counter intuitive that there is a bijective, and even biholomorphic, correspondence between spectral curves as classical systems and quantum curves as quantum systems, discovered in collaboration by the author [27, 32], prompted by the idea of Gaiotto [41]. Suppose we have a holomorphic $n$-th order ($n\geq 1$) linear ordinary differential operator $P$ globally defined on a smooth complete complex algebraic curve $C$ of genus $g(C)>1$, whose leading coefficient does not vanish anywhere on $C$. One interpretation of the above result proves that it determines a unique algebraic curve $\pi:\Sigma\longrightarrow C$, known as a spectral curve of a Higgs bundle, that recovers $P$ through the conformal limit construction of Gaiotto [41]. The passage from $P$ to $\Sigma$ includes:
@@
In terms of a local coordinate $z$ of a local neighborhood $U\subset C$, $P$ is a noncommutative polynomial in $d/dz$ of degree $n$ with coefficients in holomorphic functions in ${\mathcal{O}}_{C}(U)$. The $\hbar$-deformation changes it to a $1$-dimensional Schrödinger operator on $U$, which patches together to a globally defined operator $P^{\hbar}$. The process of semi-classical limit is equivalent to taking the “total symbol” of this differential operator. After a suitable choice of its coordinate and a conjugation action of the operator with a locally non-vanishing function, we can write
(0.4) $P=\left(\frac{d}{dz}\right)^{n}+\sum_{k=2}^{n}a_{k}(z)\left(\frac{d}{dz}\right)^{n-k},\qquad a_{k}\in{\mathcal{O}}_{C}(U)$
@@
(0.5) $\nabla_{z}:=\frac{d}{dz}+\begin{bmatrix}0&a_{2}&\cdots&a_{n-1}&a_{n}\\ -1&\\ &-1&&&\vdots\\ &&\ddots\\ &&&-1&0\end{bmatrix}\quad\text{and}\quad\Psi=\begin{bmatrix}\psi^{(n-1)}\\ \psi^{(n-2)}\\ \vdots\\ \psi^{\prime}\\ \psi\end{bmatrix}.$
Therefore, locally an $n$-th order differential equation $P\psi=0$ is always equivalent to the flatness equation $\nabla_{z}\Psi=0$ with respect to a holomorphic connection $\nabla_{z}$. In Section 3, we will translate the condition that $P$ is globally defined on $C$ into a set of properties of this connection. The requirements include $P$ acts on the line bundle $K_{C}^{-\frac{n-1}{2}}$, and the connection $\nabla$ acts on a vector bundle that has a full-flag filtration with Griffiths transversality. Such a connection is known as an oper [6, 7].
The statement we mentioned above is thus equivalent to the assertion that every oper on $C$ corresponds in a one-to-one manner to a spectral curve $\Sigma$ covering $C$. The $\hbar$-deformation of $\nabla$ is the $\hbar$-connection $\nabla^{\hbar}$ of Deligne, and the parameter $\hbar$ is identified as an element $\hbar\in H^{1}(C,K_{C})$ (see [32]), which determines the extension classes of line bundles associated with the full-flag filtration. In [27], the collaborators and the author constructed a globally defined linear differential operator on a curve $C$, or an oper, from an arbitrary spectral curve of a $G$-Hitchin system on a base curve $C$, where $G$ is a simple Lie group of adjoint type. This is a biholomorphic construction from the moduli space of spectral curves to the moduli space of opers. Therefore, it means that the spectral curve is uniquely chosen from a given oper. And because of the passage explained above, this also implies that a higher order differential operator uniquely identifies a quantization of the spectral curve, i.e., the $\hbar$-family of deformations of the starting operator, or equivalently, a Deligne’s $\hbar$-connection.
There is an important difference between holomorphic differential operators $P$ on a curve $C$ of genus $g(C)>1$, and the examples coming from enumeration problems mentioned above. When $P$, and its corresponding oper $\nabla$, is holomorphic, any solution to the equation $P\psi=0$ is everywhere holomorphic. So we do not expect it to contain any new geometric information of something that goes beyond the given context, such as topology of ¯${\overline{\mathcal{M}}}_{g,n}$ as in the examples. We expect that when we consider $P$ with irregular singularities, a whole new story begins. This is an active area of research in geometry. We refer to [2, 19, 20, 42].
@@ 0.4. A quick guide of the contents
  • The story given in Section 1.1 illustrates the model of the theory: The expansion of a solution of a differential equation ($=$ quantum curve) around its irregular singular point contains a profound amount of geometric information not obvious from the given setting.
@@
  • What we mean by a globally defined high order linear differential operator on a curve $C$ of genus $g(C)>1$ is explained in Section 3.2. There and in the following Section 3.3, we will encounter the geometric meaning of the parameter $\hbar$, how it determines the unique quantization from the starting classical spectral curve, and how the projective coordinate system appears in the global construction.

### v2

@@ Abstract.
It has been noticed since around 2007 that certain enumeration problems can be solved when an analytic or algebraic curve is identified. This curve is the key to the problem. In these lectures, a few such examples are presented. One is a detailed account on counting simple Hurwitz numbers, explaining how the problem was solved by discovering this key curve. The formula for the curve allows us to write the generating functions of Hurwitz numbers in terms of polynomials. This unexpected polynomiality produces, as a byproduct, straightforward and short proofs of the Witten-Kontsevich theorem and the $\lambda_{g}$-theorem of Faber-Pandharipande. An analogous enumeration problem associated with Catalan numbers is also presented, which has a simpler feature in terms of analysis. The asymptotic behavior of counting leads this time to the Euler characteristic of the moduli spaces of smooth curves. We then discuss another enumeration problem, the Apéry sequences. The quest of identifying the hidden curve for this case remains open.
These curves, also known as spectral curves, are discovered via solving ordinary differential equations. The counting problem of geometric origin associated with the genus 0, one marked point case is encoded in the spectral curve. It is explained that going from the $(g,n)=(0,1)$-case to arbitrary $(g,n)$ is a process of quantization of the spectral curve.
This perspective of quantization is discussed in a geometric setting, when the differential equations are linear with holomorphic coefficients, in terms of Higgs bundles, opers, and Gaiotto’s conformal limit construction. In this context, however, there are no counting problems behind the scene.
@@ 0.1. The story begins with Catalan numbers and their Laplace transform
as a basis for all solutions. It then reminds us of the quadratic formula for $z^{2}-xz+1=0$, and its possible relation to the second equation through the Weyl quantization $\begin{cases}z\longmapsto-\hbar\frac{d}{dx},\\ x\longmapsto x.\end{cases}$ Equivalently, we can find the same polynomial from the semi-classical limit of (0.2):
$\lim_{\hbar\rightarrow 0}e^{-\frac{S_{0}(x)}{\hbar}}\left(\hbar^{2}\frac{d^{2}}{dx^{2}}+\hbar x\frac{d}{dx}+1\right)e^{\frac{S_{0}(x)}{\hbar}}=\big(S_{0}(x)^{\prime}\big)^{2}+xS_{0}(x)^{\prime}+1,\qquad z=-S_{0}(x)^{\prime}.$
@@
The curve in common appears to be $x=z+\frac{1}{z}.$ Is there anything significant here? We will see that this curve tells us a rich story of geometry that is not obvious from the shape of these differential equations. As explained in the main text (1.2), $z(x)$ is an unconventional generating function of Catalan numbers $C_{m}:1,2,5,14,42,132,429,1430,4862,16796\dots$ One interpretation of these numbers is the count of cell-decompositions of a two-dimensional sphere $S^{2}$ with one 0-cell and $m$ 11-cells, while not allowing rotation symmetry to avoid complication coming from automorphism count (see Remark 2.29).
Figure 0.2. A cell-decomposition of a curve of genus 0 with one 0-cell. The arrow on a half-edge indicates no rotation symmetry is allowed.
@@
What do these geometric quantities of moduli spaces have anything to do with the differential equation (0.2) above? We will see that its solution $\psi(x,\hbar)$ is the generating function of these generating functions $F_{g,n}$ for all $g\geq 0$ and $n>0$. In this context, since the curve $x=z+1/z$ is the semi-classical limit of (0.2), the differential equation is called the quantum curve [1, 23, 53] associated with the classical curve $x=z+1/z$. The Catalan numbers are the count of cell-decompositions of the $(g,n)=(0,1)$ geometry. Hence going from $z(x)$ to $F_{g,n}(t_{1},\dots,t_{n})$ for arbitrary $(g,n)$ is the process of quantization, in analogy of counting graphs of higher genera as Feynman diagrams representing interactions.
At the same time, the fact that generating functions $F_{g,n}(t_{1},\dots,t_{n})$ know the $\psi$-class intersection numbers of ${\overline{\mathcal{M}}}_{g,n}$ indicates that they give the Gromov-Witten invariants $GW_{g,n}(\bullet)$ of a point. Therefore, from the point of view of Mariño [69] and Bouchard-Klemm-Mariño-Pasquetti [14], we understand that the spectral curve $x=z+1/z$ is the mirror $B$-model, corresponding to the $A$-model on a trivial symplectic manifold, i.e., a point. From the Catalan numbers to this point of view of mirror symmetry is the theme of Section 1.1.
The polynomiality of some quantities appearing in enumerative problems associated with the moduli spaces ${\overline{\mathcal{M}}}_{g,n}$ was discovered in an earlier paper [80], which was the key to the solution [37] of the Hurwitz number conjecture of Bouchard and Mariño [15]. The polynomiality of the generating functions of simple Hurwitz numbers established in [80] presented another surprise: it gives simple few-line proofs of Witten’s conjecture [98] and the $\lambda_{g}$-conjecture of Faber-Pandharipande [39, 40].
The stories coming out of Hurwitz numbers are weaved in Section 2. The hidden curve in this context is the Lambert curve (2.15), whose role in discovering the polynomiality is explained in detail. Bouchard and Mariño [15] identified the Lambert curve through the limiting process of the mirror curves of toric Calabi-Yau $3$-folds [14], and using Lagrange’s Inversion Formula [97]. Appendix gives an elementary analysis that leads to this identification of the Lambert curve. In Section 2.6, we will show that the Lambert curve is a simple consequence of tree counting, and the formula for the curve, i.e., the Lambert function, directly follows from a combinatorial identity of trees.
@@
    $z(x)=\sum_{m=0}^{\infty}\frac{C_{m}}{x^{2m+1}},\hskip 23.49976ptx=z+\frac{1}{z}$
@@
  • This effect is explained in Section 2 using Hurwitz numbers. The starting sequence is the number of rooted trees on $k$ nodes and their generating function,
@@
  • In Section 2.3 it is explained that the Laplace transform of the $(g,n)$-Hurwitz number in the stable region 2g2+n>02g-2+n>0 is a polynomial in the $t$-variables. We will prove that the Laplace transform of the combinatorial formula, the cut-and-join equation of [48, 94], is a differential equation, and that it automatically proves the Witten-Kontsevich theorem on the $\psi$-class intersection numbers and the Faber-Pandharipande theorem on the $\lambda_{g}$-conjecture [80], through the ELSV formula [36].
@@ 0.2. A miraculous integer sequence
$\widetilde{GW}_{0,1}(n):=(-1)^{n}(n!)^{2}\sum_{\begin{subarray}{c}\ell+m=n\\ \ell,m\geq 0\end{subarray}}\frac{(2\ell+m)!(\ell+2m)!}{(\ell!)^{5}(m!)^{5}}\bigg(1+(m-\ell)\big(H_{2\ell+m}+2H_{\ell+2m}-5H_{m}\big)\bigg)$
@@
Then we have (see [19, 47, 99]):
(0.3) $A_{n}=\widetilde{GW}_{0,1}(n)\hskip 23.49976pt\text{for all }\;n.$
The first few terms are: $1,5,73,1445,33001,819005,21460825,584307365,16367912425,\dots$. What is unbelievable about this equality is that $\widetilde{GW}_{0,1}(n)$ as defined is a positive integer, and has a simple expression as in the second line. What we know about (0.3) so far includes the following:
  • The formula comes from the mirror symmetry, a geometric relation between the Gromov-Witten invariants of a complex $3$-dimensional variety and the Gauss-Manin connection associated with its mirror partner. In this particular context, we refer to Golyshev-Zagier [47] and Zagier [99] for an inspiring account of this interplay.
  • The quantity $\widetilde{GW}_{0,1}(n)$, the quantum period, was obtained by Coates, Corti, Galkin, and Kasprzyk [19], which is essentially the same as the genus 0, 11-marked point case of the degree $n$ Gromov-Witten invariant of a Fano $3$-fold known as $V_{12}$. The generating function of these numbers, after appropriate modifications including the Borel-Laplace transform, satisfies a Quantum Differential Equation. A brief background will be explained in Section 4.
@@ 0.3. The unique quantization for the holomorphic cases
It is therefore rather counter intuitive that there is a bijective, and even biholomorphic, correspondence between spectral curves as classical systems and quantum curves as quantum systems, discovered in [28, 33], prompted by the idea of Gaiotto [42]. Suppose we have a holomorphic $n$-th order ($n\geq 1$) linear ordinary differential operator $P$ globally defined on a smooth complete complex algebraic curve $C$ of genus $g(C)>1$, whose leading coefficient does not vanish anywhere on $C$. One interpretation of the above result proves that it determines a unique algebraic curve $\pi:\Sigma\longrightarrow C$, known as a spectral curve of a Higgs bundle, that recovers $P$ through the conformal limit construction of Gaiotto [42]. The passage from $P$ to $\Sigma$ includes:
@@
In terms of a local coordinate $z$ of a local neighborhood $U\subset C$, $P$ is a noncommutative polynomial in $d/dz$ of degree $n$ with coefficients in holomorphic functions in ${\mathcal{O}}_{C}(U)$. The $\hbar$-deformation changes it to a $1$-dimensional Schrödinger operator on $U$, which patches together to a globally defined operator $P^{\hbar}$ on $C$. The process of semi-classical limit is equivalent to taking the “total symbol” of this differential operator. After a suitable choice of a local coordinate and a conjugation action of the operator with a locally non-vanishing function, we can write
(0.4) $P=\left(\frac{d}{dz}\right)^{n}+\sum_{k=2}^{n}a_{k}(z)\left(\frac{d}{dz}\right)^{n-k},\hskip 23.49976pta_{k}\in{\mathcal{O}}_{C}(U)$
@@
(0.5) $\nabla_{z}:=\frac{d}{dz}+\begin{bmatrix}0&a_{2}&\cdots&a_{n-1}&a_{n}\\ -1&\\ &-1&&&\vdots\\ &&\ddots\\ &&&-1&0\end{bmatrix}\qquad\text{and}\qquad\Psi=\begin{bmatrix}\psi^{(n-1)}\\ \psi^{(n-2)}\\ \vdots\\ \psi^{\prime}\\ \psi\end{bmatrix}$.
Therefore, locally an $n$-th order differential equation $P\psi=0$ is always equivalent to the flatness equation $\nabla_{z}\Psi=0$ with respect to a holomorphic connection $\nabla_{z}$. In Section 3, we will translate the condition that $P$ is globally defined on $C$ into a set of properties of this connection. The requirements include that $P$ acts on the line bundle $K_{C}^{-\frac{n-1}{2}}$, and that the connection $\nabla$ acts on a vector bundle that has a full-flag filtration with Griffiths transversality. Such a connection is known as an oper [7, 8].
The statement we mentioned above is thus equivalent to the assertion that every oper on $C$ corresponds in a one-to-one manner to a spectral curve $\Sigma$ covering $C$. The $\hbar$-deformation of $\nabla$ is the $\hbar$-connection $\nabla^{\hbar}$ of Deligne, and the parameter $\hbar$ is identified as an element $\hbar\in H^{1}(C,K_{C})$ (see [33]), which determines the extension classes of line bundles associated with the full-flag filtration. In [28], a globally defined linear differential operator on a curve $C$, or an oper, is constructed from an arbitrary spectral curve of a $G$-Hitchin system on a base curve $C$, where $G$ is a simple Lie group of adjoint type. This gives a biholomorphic map from the moduli space of spectral curves to the moduli space of opers. Bijection means that the original spectral curve is uniquely determined by the corresponding oper. And the construction also implies that a higher order differential operator uniquely identifies a quantization of the spectral curve, i.e., the $\hbar$-family of deformations of the starting operator, or equivalently, a Deligne’s $\hbar$-connection.
There is an important difference between holomorphic differential operators $P$ on a curve $C$ of genus $g(C)>1$, and the examples coming from enumeration problems mentioned above. When $P$, and its corresponding oper $\nabla$, is holomorphic, any solution to the equation $P\psi=0$ is everywhere holomorphic. So we do not expect it to contain any new geometric information of something that goes beyond the given context, such as topology of ${\overline{\mathcal{M}}}_{g,n}$ as in the examples. We expect that when we consider $P$ with irregular singularities, a whole new story begins. This is an active area of research in geometry. We refer to [2, 20, 21, 43].
@@ 0.4. A quick guide of the contents
  • The story given in Section 1.1 illustrates the model of the theory: The expansion of a solution of a differential equation (== quantum curve) around its irregular singular point contains a profound amount of geometric information not obvious from the given setting.
@@
  • What we mean by a globally defined high order linear differential operator on a curve $C$ of genus $g(C)>1$ is explained in Section 3.2. There and in the following Section 3.3, we will encounter the geometric meaning of the parameter $\hbar$, how it determines the unique quantization from the starting classical spectral curve, and how the projective coordinate system [49] appears in the global construction.